OpenStudy (anonymous):

using the following equation, find the center and radius of the circle. x^2 + 4x + y^2 - 6y= -4

OpenStudy (anonymous):

Please help I have none of my formulas!

OpenStudy (lucaz):

you have to complete the squares to get the equation of a circle in standard form

OpenStudy (anonymous):

What? How can I complete the squares when it's y and x

OpenStudy (lucaz):

you have to complete the square with x and y, then you wil get: (x-a)^2 + (y-b)^2 = r^2 x-a is the change in x y-b is the change in y r is the radius

OpenStudy (anonymous):

I'm sorry I'm just trying to understand that. So how to I get the radius, r, if that's all letters? I must be being very thick right now I'm sorry

OpenStudy (anonymous):

.?

OpenStudy (whpalmer4):

Okay, you've got x^2 + 4x + y^2 - 6y = -4 you need to get (x-a)^2 + (y-b)^2 = r^2 x^2 + 4x + <some number> is a perfect square if you can find a number (we'll call it c) such that (x+c)^2 = (x+c)(x+c) = x^2 + 4x + <some number>

OpenStudy (whpalmer4):

(x+c)(x+c) = x^2 + cx + cx + c^2 = x^2 + 2cx + c^2 if we can find a value of c so that 4x = 2cx, then we can write x^2 + 4x + <something> = (x+c)^2 and that will be completing the square on x. then we'll do the same procedure to complete the square on y.

OpenStudy (anonymous):

I seriously have no idea. Do I have to find the number that x^2+ 4x equals?

OpenStudy (whpalmer4):

4x = 2cx can you solve that for c?

OpenStudy (whpalmer4):

hint: you can divide both sides by x, to get 4x/x = 2cx/x or 4 = 2c

OpenStudy (anonymous):

Then 4=2c /2. /2 2=c

OpenStudy (whpalmer4):

right! so (x+c)^2 = (x+2)^2 = (x+2)(x+2) = x^2 + 2x + 2x + 4 = x^2 + 4x + 4 well, that's just like the x stuff we have in our original equation, right? we just need to find a spare "+4" somewhere :-)

OpenStudy (whpalmer4):

we had x^2 + 4x + y^2 - 6y= -4 we'd like to rewrite the x^2 + 4x part as (x+c)^2 = (x+2)^2 but to do that we need a "+4" on the left side of the equation. we can just make one! 4 - 4 = 0, right? so we can insert x^2 + 4x + 4 - 4 + y^2 - 6y = -4 we haven't changed the value of the left side or the right side by doing that. [x^2 + 4x + 4] - 4 + y^2 - 6y = -4 now we can rewrite the stuff in the brackets as (x+2)^2 (x+2)^2 - 4 + y^2 - 6y = -4 progress! now we need to do the same process with y^2 - 6y

OpenStudy (whpalmer4):

does that make sense so far?

OpenStudy (whpalmer4):

if y^2 - 6y + <something> = y^2 + 2cy + c^2 -6y = 2cy -6 = 2c c = -3 so we can rewrite y^2 - 6y + 9 as (y-3)^2

OpenStudy (whpalmer4):

we just need to find (or create) a +9 in our equation. (x+2)^2 - 4 + y^2 - 6y = -4 (x+2)^2 - 4 + y^2 - 6y + 9 - 9 = -4 (x+2)^2 -4 + [y^2 - 6y + 9] - 9 = -4 replace the stuff in brackets by (y-3)^2 (x+2)^2 - 4 + (y-3)^2 - 9 = -4 now we just move the remaining numbers to the right hand side (x+2)^2 -4 + 4 + (y-3)^2 - 9 + 9 = - 4 + 4 + 9 (x+2)^2 + (y-3)^2 = 9 that's the form we want: (x-h)^2 + (y-k)^2 = r^2 to make those match, we have (x+2) = (x-h) 2 = -h h = -2 (y-3) = (y-k) -3 = -k k = 3 r^2 = 9 r = 3

OpenStudy (whpalmer4):

the general form of the circle equation: (x-h)^2 + (y-k)^2 = r^2 gives us the center at (h,k) and the radius is r if our equation is (x+2)^2 + (y-3)^2 = 9 and h = -2, k = 3, and r = 3 what is the radius? what is the center?

OpenStudy (whpalmer4):

radius is 3, because r = 3 center is at (-2, 3) because center is at (h, k) and h = -2, k = 3

OpenStudy (anonymous):

You are very nice thanks for explaining that all! Honestly I didn't study at all because my end date is today and I have to finish this so I have no idea what I'm doing so thank you!

OpenStudy (whpalmer4):

I offer a full money-back guarantee. If tomorrow this still doesn't make sense, tag me on a question and I'll try explaining it again until it does :-)

OpenStudy (anonymous):

Haha ok thank you